THE Find sino 13 9 / 7 b) tangent 0 = ?Exact Values except when the calc. is stipulated. Can you please explain how you solve this with all work and clear handwriting.
Answer #1

Copposite) trypotenuse sa Using phythagorean identity, at +8=133 22-132-22=169-49=120 a Vizo = 2130 151 (adjacent sind opposi